### v1

@@ Abstract
Compact neural networks are specially designed for applications on edge devices with faster inference speed yet modest performance. However, training strategies of compact models are borrowed from that of conventional models at present, which ignores their difference in model capacity and thus may impede the performance of compact models. In this paper, by systematically investigating the impact of different training ingredients, we introduce a strong training strategy for compact models. We find that the appropriate designs of re-parameterization and knowledge distillation are crucial for training high-performance compact models, while some commonly used data augmentations for training conventional models, such as Mixup and CutMix, lead to worse performance. Our experiments on ImageNet-1K dataset demonstrate that our specialized training strategy for compact models is applicable to various architectures, including GhostNetV2, MobileNetV2 and ShuffleNetV2. Specifically, equipped with our strategy, GhostNetV3 1.3$\times$ achieves a top-1 accuracy of 79.1% with only 269M FLOPs and a latency of 14.46ms on mobile devices, surpassing its ordinarily trained counterpart by a large margin. Moreover, our observation can also be extended to object detection scenarios. PyTorch code and checkpoints can be found at https://github.com/huawei-noah/Efficient-AI-Backbones.
@@ 1 Introduction
Based on our investigation, we develop a specialized training recipe for compact models. Experiments on ImageNet-1K dataset verify the superiority of our proposed recipe. Specifically, the GhostNetV2 model trained with our recipe significant outperforms that trained with the previous strategy in terms of both top-1 accuracy and latency (Figure 1). Experiments on other efficient architectures such as MobileNetV2 and ShuffleNetV2 further confirm the generalizability of the proposed recipe.
@@ 4.1 Re-parameterization
Re-parameterization has proved its effectiveness in conventional convolutional models [7, 6]. Inspired by their success, we introduce re-parameterization into compact models by adding repetitive branches equipped with BatchNorm layers. Our design of re-parameterization GhostNetV2 is presented in Figure 2(b). It is worth noting that we introduce a 1$\times$1 depth-wise convolution branch into re-parameterized 3$\times$3 depth-wise convolution. Experimental results confirm its positive effect on the performance of compact models. Furthermore, the experiments thoroughly explore the optimal number of repetitive branches.
@@ 5.1 Re-parameterization
To better comprehend the advantages of integrating re-parameterization into the training of compact models, we perform an ablation study to assess the impact of re-parameterization on different sizes of GhostNetV2. The results are presented in Table 2. The adoption of re-parameterization, while keeping other training settings unchanged, results in a significant improvement in performance compared to directly training the original GhostNetV2 models.
Table 1: Top-1 accuracy of different versions of GhostNetV2 models trained with or without re-parameterization.
@@ 5.2 Knowledge distillation
We also explore the impact of combining re-parameterization and knowledge distillation on the performance of GhostNetV2. As shown in Table 6, the results indicate a significant improvement in performance (up to 79.13%) due to the utilization of knowledge distillation. Moreover, it emphasizes the importance of the 1$\times$1 depth-wise convolution in re-parameterization. These findings underscore the importance of investigating various techniques and their potential combinations to enhance the performance of compact models.
@@ 5.3 Learning schedule
Figure 3: The top-1 validation accuracy for various learning rates of GhostNetV2.
@@
The effect of weight decay on the top-1 accuracy of GhostNetV2 is shown in Table 7. The results indicate that a large weight decay significantly diminishes the model’s performance. Therefore, we have retained a weight decay value of 0.05 for GhostNetV2, given its effectiveness for compact models.
@@ 5.4 Data Augmentation
To compare the impact of different data augmentation schemes on the performance of lightweight models, we train cnn-based GhostNetV2 and ViT-based DeiT-tiny models under distinct augmentation strategies. The results are presented in Table 8. It is observed that random augmentation and random erasing are advantageous for both GhostNetV2 and DeiT-tiny. Conversely, Mixup and CutMix have a detrimental effect, which are then considered unsuitable for compact models.
Table 8: comparison results of different combinations of data augmentation schemes on GhostNetV2 and DeiT-tiny.
AutoAug RandAug Mixup CutMix Erasing GhostNetV2 DeiT-tiny
@@ 5.5 Comparison with other compact models
Compared with MobileNetV2 [21], GhostNetV2 1.0$\times$ achieves a 5.1% improvement while maintaining almost the same latency (7.81ms vs. 7.96ms). GhostNetV2 1.3$\times$ also exhibits improved top-1 accuracy compared to MobileNeXt and EfficientNet-B0, which are 3.0% and 2.8%, respectively. In particular, when compared with the powerful manually designed MobileOne models, GhostNetV3 1.0$\times$ outperforms MobileOne-S1 in terms of top-1 accuracy by 1.2% with only half the latency required. GhostNetV3 1.3$\times$ also outperforms MobileOne-S2 by 1.7% while costing only 60% of the latency. Moreover, when GhostNet 1.6$\times$ achieves a higher top-1 accuracy than MobileOne-S4 (80.4% vs. 79.4%), MobileOne’s latency is 2.8$\times$ slower than that of GhostNetV3 on CPU.
@@
Figure 5 presents a comprehensive performance comparison of various compact models. The left and right figures illustrate the FLOPs and the latency measured on a mobile phone, respectively. Notably, our trained GhostNetV2 stands out by exhibiting the best balance between latency and top-1 accuracy on mobile devices.

### v2

@@ Abstract
Compact neural networks are specially designed for applications on edge devices with faster inference speed yet modest performance. However, training strategies of compact models are borrowed from that of conventional models at present, which ignores their difference in model capacity and thus may impede the performance of compact models. In this paper, by systematically investigating the impact of different training ingredients, we introduce a strong training strategy for compact models. We find that the appropriate designs of re-parameterization and knowledge distillation are crucial for training high-performance compact models, while some commonly used data augmentations for training conventional models, such as Mixup and CutMix, lead to worse performance. Our experiments on ImageNet-1K dataset demonstrate that our specialized training strategy for compact models is applicable to various architectures, including GhostNetV3, MobileNetV2 and ShuffleNetV2. Specifically, equipped with our strategy, GhostNetV3 1.3$\times$ achieves a top-1 accuracy of 79.1% with only 269M FLOPs and a latency of 14.46ms on mobile devices, surpassing its ordinarily trained counterpart by a large margin. Moreover, our observation can also be extended to object detection scenarios. PyTorch code and checkpoints can be found at https://github.com/huawei-noah/Efficient-AI-Backbones.
@@ 1 Introduction
Based on our investigation, we develop a specialized training recipe for compact models. Experiments on ImageNet-1K dataset verify the superiority of our proposed recipe. Specifically, the GhostNetV3 model trained with our recipe significant outperforms that trained with the previous strategy in terms of both top-1 accuracy and latency (Figure 1). Experiments on other efficient architectures such as MobileNetV2 and ShuffleNetV2 further confirm the generalizability of the proposed recipe.
@@ 4.1 Re-parameterization
Re-parameterization has proved its effectiveness in conventional convolutional models [7, 6]. Inspired by their success, we introduce re-parameterization into compact models by adding repetitive branches equipped with BatchNorm layers. Our design of re-parameterization GhostNetV3 is presented in Figure 2(b). It is worth noting that we introduce a 1$\times$1 depth-wise convolution branch into re-parameterized 3$\times$3 depth-wise convolution. Experimental results confirm its positive effect on the performance of compact models. Furthermore, the experiments thoroughly explore the optimal number of repetitive branches.
@@ 5.1 Re-parameterization
To better comprehend the advantages of integrating re-parameterization into the training of compact models, we perform an ablation study to assess the impact of re-parameterization on different sizes of GhostNetV3. The results are presented in Table 2. The adoption of re-parameterization, while keeping other training settings unchanged, results in a significant improvement in performance compared to directly training the original GhostNetV3 models.
Table 1: Top-1 accuracy of different versions of GhostNetV3 models trained with or without re-parameterization.
@@ 5.2 Knowledge distillation
We also explore the impact of combining re-parameterization and knowledge distillation on the performance of GhostNetV3. As shown in Table 6, the results indicate a significant improvement in performance (up to 79.13%) due to the utilization of knowledge distillation. Moreover, it emphasizes the importance of the 1$\times$1 depth-wise convolution in re-parameterization. These findings underscore the importance of investigating various techniques and their potential combinations to enhance the performance of compact models.
@@ 5.3 Learning schedule
Figure 3: The top-1 validation accuracy for various learning rates of GhostNetV3.
@@
The effect of weight decay on the top-1 accuracy of GhostNetV3 is shown in Table 7. The results indicate that a large weight decay significantly diminishes the model’s performance. Therefore, we have retained a weight decay value of 0.05 for GhostNetV3, given its effectiveness for compact models.
@@ 5.4 Data Augmentation
To compare the impact of different data augmentation schemes on the performance of lightweight models, we train cnn-based GhostNetV3 and ViT-based DeiT-tiny models under distinct augmentation strategies. The results are presented in Table 8. It is observed that random augmentation and random erasing are advantageous for both GhostNetV3 and DeiT-tiny. Conversely, Mixup and CutMix have a detrimental effect, which are then considered unsuitable for compact models.
Table 8: comparison results of different combinations of data augmentation schemes on GhostNetV3 and DeiT-tiny.
AutoAug RandAug Mixup CutMix Erasing GhostNetV3 DeiT-tiny
@@ 5.5 Comparison with other compact models
Compared with MobileNetV2 [21], GhostNetV3 1.0$\times$ achieves a 5.1% improvement while maintaining almost the same latency (7.81ms vs. 7.96ms). GhostNetV3 1.3$\times$ also exhibits improved top-1 accuracy compared to MobileNeXt and EfficientNet-B0, which are 3.0% and 2.8%, respectively. In particular, when compared with the powerful manually designed MobileOne models, GhostNetV3 1.0$\times$ outperforms MobileOne-S1 in terms of top-1 accuracy by 1.2% with only half the latency required. GhostNetV3 1.3$\times$ also outperforms MobileOne-S2 by 1.7% while costing only 60% of the latency. Moreover, when GhostNet 1.6$\times$ achieves a higher top-1 accuracy than MobileOne-S4 (80.4% vs. 79.4%), MobileOne’s latency is 2.8$\times$ slower than that of GhostNetV3 on CPU.
@@
Figure 5 presents a comprehensive performance comparison of various compact models. The left and right figures illustrate the FLOPs and the latency measured on a mobile phone, respectively. Notably, our trained GhostNetV3 stands out by exhibiting the best balance between latency and top-1 accuracy on mobile devices.